# Cross-lingual Offensive Language Detection: A Systematic Review of Datasets, Transfer Approaches and Challenges

### v1

@@ 6.1.1. Label-Based Transfer
It involves leveraging parallel corpora to transfer annotations from a source language to a target language. Parallel corpora consist of sentence-aligned source and target language data, ensuring higher quality than machine-generated translations. Since data in the source language is paired with target language data, labels in the source language can be directly projected onto the target language. For example, if a sentence in the source language is labelled as hateful, its corresponding text in the target language can be inferred to have the same label. It can effectively create labelled data for the target language without manual annotation, especially useful when building datasets for languages where annotating from scratch might be challenging due to linguistic complexities or lack of expert annotators.
@@ 6.2. Feature Transfer
Feature transfer approaches transform linguistic features to aid in the detection of offensive language. Rather than translating source texts or projecting source labels, they focus on extracting salient linguistic features from source and target languages, and aligning them into a shared feature space, retaining the essence of the text consistent across languages.
Cross-lingual Word Embeddings (CLWEs), also known as Multilingual Word Embeddings (MWEs), are commonly used for feature transfer in offensive language detection. They train monolingual word embeddings (such as Word2Vec or FastText) on multiple languages, yielding vectors that capture semantic similarities relying on shared representations between languages and thus enabling representations that are common across languages.
In cross-lingual offensive language detection, various approaches utilise existing CLWEs or retrofit CLWEs. We discuss three main techniques: multilingual distributional representations, multilingual contextualised representations, and retrofitting word embeddings.
@@ 6.2.1. Multilingual Distributional Representations
Pre-trained distributional word embeddings are most commonly used as embedding layers in neural networks for different language data inputs. Among these, MUSE embeddings have emerged as a predominant choice. They are extensively utilised to extract multilingual features and are often integrated with both traditional machine learning models, such as Gradient Boosted Decision Trees (Arango et al., 2022), and advanced deep learning architectures (i.e. LSTM (Pamungkas and Patti, 2019; Pamungkas et al., 2020, 2021a; Arango et al., 2021; Burtenshaw and Kestemont, 2021), BiLSTM (Bigoulaeva et al., 2021, 2022, 2023; Jiang and Zubiaga, 2021), CNN-GRU (Deshpande et al., 2022; Aluru et al., 2021, 2020), Capsule Networks (Jiang and Zubiaga, 2021), Sluice networks (Ousidhoum et al., 2019)), and BERT (Pamungkas et al., 2020)). Other notable multilingual distributional vectors, such as Multilingual GloVe and FastText, have also been employed in architectures like BiLSTM (Chiril et al., 2019) and Capsule Networks (Jiang and Zubiaga, 2021). In addition, Babylon multilingual embeddings have been paired with Sluice networks for multilingual hate speech detection (Ousidhoum et al., 2019). Bansal et al. (2020) integrate bilingual switching features into the Hierarchical Attention Network (HAN) architecture, enhancing the transfer knowledge for cross-lingual detection.
@@ 6.2.3. Retrofitting Word Embeddings
Retrofitting pre-trained word embeddings by infusing multilingual domain knowledge stands as a promising strategy to amplify the semantic relationships across languages, thereby enhancing the efficacy and scalability of models in cross-lingual offensive language detection. Pant and Dadu (2020) employ a supervised FastText model trained on the sarcasm detection corpus (Swami et al., 2018) to improve the identification of hate speech content in English and Hindi languages. Arango et al. (2021) construct hate speech-specific word embeddings by aligning monolingual Word2Vec embeddings using hate-specific bilingual dictionaries (such as HurtLex), which are able to capture non-traditional translations of words between languages. Additionally, Hahn et al. (2021) learn semantic subspace-based representations to model profane languages on both word and sentence levels. Jiang and Zubiaga (2023) propose a cross-lingual domain-aware semantic specialisation system to construct sexism-aware word embeddings. They retrofit pre-trained FastText word vectors by integrating in-domain and out-of-domain linguistic knowledge (such as lexico-semantic relations) into the specialised feature space.
@@ 6.2.4. Infusing Additional Features
Additional external features are commonly used in identifying hate speech across languages, such as domain-specific (Pamungkas and Patti, 2019; Hahn et al., 2021), language-specific (Zhou et al., 2023; Arango et al., 2021), and typographic features (Kumar and Sachdeva, 2022; Markov et al., 2021). Specifically, domain-specific features are essential to assist the hate-related knowledge transfer. The multilingual lexicon, HurtLex, has been employed in multiple studies (Pamungkas et al., 2020, 2021a; Pamungkas and Patti, 2019; Jiang and Zubiaga, 2021). Some researchers also utilise hateful word pairs to construct domain-aware or semantic-based representations across languages (Jiang and Zubiaga, 2023; Hahn et al., 2021; Arango et al., 2021), due to the lack of hate-specific patterns in general-purpose multilingual embeddings. In addition, infusing language-specific features can help integrate linguistic and cultural knowledge for geographically sensitive tasks. It can be in different forms, such as language switching pattern matrix (Bansal et al., 2020), bilingual pairs (Jiang and Zubiaga, 2023), cross-cultural similarities (Arango et al., 2021), and social dynamics among users (Nozza, 2021). This feature can provide a richer understanding of cultural dimensions in hateful content (Zhou et al., 2023). Furthermore, typographic features provide a language-agnostic and domain-agnostic perspective to help identify online hate due to their shared meanings across languages, which includes capitals, punctuations, and emojis (Kumar and Sachdeva, 2022; Markov et al., 2021; Corazza et al., 2020). Stylometric features also remain persistent with respect to language variations, as the writing style of toxic content can be correlated to emotional profile of social media users (Markov et al., 2021).
@@ 6.3. Parameter Transfer
Performance of NLP models is dependent on the values of model parameters. Parameter transfer assumes that some parameters or prior distributions of hyperparameters are transferred between different languages within one model or between individual models. Most parameter transfer approaches rely on inductive transfer learning in cross-lingual offensive language detection (Pan and Yang, 2010). Inspired by Pikuliak et al. (2021), we outline three distinct scenarios for all surveyed studies, depending on how the source and target data are utilised during the process of parameter transfer, namely zero-shot transfer, joint learning, and cascade learning (see Figure 5). Only source data is used for training in zero-shot transfer, while both source and target data are applied during the training process in joint learning and cascade learning. Joint learning utilises both simultaneously to train the model, but cascade learning leverages them in the different stages of training.
@@
Multilingual PLMs |
Domain Adaptation |
Multi-source Training |
Multi-channel Learning |
Multi-task Learning |
Meta Learning |
Ensemble Learning |
@@ 6.3.1. Transfer Scenarios
Zero-Shot Transfer
It is a promising scenario for cross-lingual offensive language detection when labelled data in the target language is scarce or lacking, by enabling detection in a language not seen during training. The model is trained on labelled source data from one or more languages not including the target language, and is then tested on the unseen target language. This is sometimes called direct or model transfer. We found 25 papers under this transfer paradigm, including subtypes:
Single Source to Single Target: Considered a type of domain adaptation (see Section 6.3.2), models are trained on labelled data from a single source language, then used on an unseen target language (Glavaš et al., 2020; Firmino et al., 2021; Tiţa and Zubiaga, 2021; Rodríguez et al., 2021; Pelicon et al., 2021a; Nozza, 2021; Montariol et al., 2022; Eronen et al., 2022; Zhou et al., 2023; Gupta et al., 2022; Ranasinghe and Zampieri, 2021a, b; Gaikwad et al., 2021; Muti and Barrón-Cedeño, 2022; Das et al., 2022; Plaza-del arco et al., 2023). It is the most frequently observed (62.5%) of zero-shot transfer strategies in our survey.
Multiple Sources to Single Target: Models are trained on labelled data from multiple source languages (excluding the target language) and then evaluated on a single target language (Ryzhova et al., 2022; Montariol et al., 2022; Aluru et al., 2020). It is a type of multi-source training strategy where the combination of training data excludes the target (see Section 6.3.2).
@@
It is worth highlighting that these methods rely on multilingual PLMs, especially often used for Single Source to Single Target and Multiple Sources to Single Target scenarios. mBERT and XLM-R are two of the most prominent multilingual PLMs in surveyed papers, because of their popularity, widespread adoption and SOTA performance across multilingual NLP tasks. They together account for 80.8% of papers in the zero-shot transfer scenario. Additionally, some papers release domain- and language-specific PLMs in a multilingual setting (Gupta et al., 2022), choose other multilingual PLMs like XLM (Stappen et al., 2020) and language-specific MuRIL (Gupta et al., 2022; Das et al., 2022), or combine multilingual word embeddings with monolingual neural models like LSTM (Pamungkas and Patti, 2019; Jiang and Zubiaga, 2021) and BERT (Pamungkas et al., 2020).
Joint Learning
It typically involves training a model between languages based on multiple channels, tasks or objectives simultaneously, or jointly assimilating multiple types of data or modalities. The model might have some parameters that are shared across languages as well as some that are specific to each, so it is also called parameter sharing (Pikuliak et al., 2021). The transfer of knowledge happens only on shared parameters. Changes to the shared parameters can affect each other between languages. They are interrelated and can provide complementary information.
@@
Cascade Learning
In the cascade learning scenario, both source and target data are used at different times during the training. First, an existing PLM is directly used or a new model is trained with source data, and then the trained model is fine-tuned with target data, where subsets of parameters are shared between different stages of training and fine-tuning. Different strategies for training and fine-tuning have been used in 19 of our surveyed papers.
Many use existing SOTA pre-trained multilingual models serving as the foundation,171717Cascade learning fine-tunes PLMs using target data, while zero-shot transfer fine-tunes PLMs using non-target data. while some studies train a model with one or multiple source datasets. Subsequently, the fine-tuning stage in cascade learning is applied across an entire (Kapoor et al., 2019; Firmino et al., 2021; Aluru et al., 2021; Pelicon et al., 2021b; Husain and Uzuner, 2022) or only a limited subset of target data (i.e. few-shot learning) (Ranasinghe and Zampieri, 2021a; Pelicon et al., 2021b), optimising the model adaptability. R”̈ottger et al. (2022a) indicate that an initial training phase on source data (English) could increase model performance when there is little fine-tuning data in the target language, where source data can partly substitute target data in few-shot settings. Some researchers delve into consistent fine-tuning (Ryzhova et al., 2022; Mozafari et al., 2022; Awal et al., 2023) or iterative fine-tuning (i.e. meta learning) (Mozafari et al., 2022; Awal et al., 2023), performing multiple rounds of fine-tuning on the PLMs to further enhance its ability to transfer knowledge.
Additionally, given the limited availability of target data, data augmentation techniques have emerged. They aim to enrich the target dataset to enable fine-tuning, by enhancing label diversity or expanding the text corpus. Some studies have attempted to generate pseudo labels for unlabelled target datasets (Bigoulaeva et al., 2021, 2022, 2023; Zia et al., 2022) (see Section 6.1.1), while others augment and refine pseudo texts by exploring advanced techniques, such as word alignment to project samples (Shi et al., 2022), adversarial algorithm to generate attack samples (Ryzhova et al., 2022), and domain-specific target data filtering (Glavaš et al., 2020).
@@ 6.3.2. Hybrid Transfer Strategies
According to the utilisation of source and target data, these strategies can be broadly covered in one or more parameter transfer scenarios mentioned in Section 6.3.1. That is, these strategies are not mutually exclusive. A single scenario can seamlessly integrate multiple transfer strategies (see Table 6).
Multilingual Pre-Trained Language Models
@@
These models are regarded as a cornerstone of parameter-level transfer techniques for low-resource scenarios, especially in cross-lingual offensive language detection. They are trained on large-scale multilingual corpora, learning commonalities and differences between languages. Therefore, a single PLM can share the same underlying sub-word vocabulary and semantic representations across languages (Devlin et al., 2019). Multilingual PLMs can be initialised as the base model with pre-trained parameters, then fine-tuned on available hate speech data in all three scenarios of parameter transfer (zero-shot transfer, joint learning and cascade learning) (Muti and Barrón-Cedeño, 2022; Firmino et al., 2021). Beyond this, they can also employed in the other two transfer levels (data and feature). They are able to augment target datasets by predicting pseudo labels for unlabelled data and generating synthetic target texts. And they can be used as multilingual representations, where subsequent models are constructed based on these source representations (Rodríguez et al., 2021).
Domain Adaptation
@@
It involves taking a model trained (or pre-trained) in the source language and adapting its parameters to initialise a new model for the low-resource language, aiming to leverage source data to improve performance on target data (Lee et al., 2021; Glavaš et al., 2020). Its advantage is that it can capture offensive language knowledge and linguistic similarities between the languages within the model’s parameters. In the zero-shot transfer scenario, parameters of the trained model are entirely or partially used on the target data for prediction (Gaikwad et al., 2021; Ranasinghe and Zampieri, 2021c). In addition, cascade learning adopts a more subtle approach. That is, the entire or specific subsets of the model’s parameters (like specific layers) are frozen to retain their original state (Ranasinghe and Zampieri, 2021a), while others are fine-tuned to better align with the target data, often using a smaller target dataset (as few-shot learning) (Pelicon et al., 2021b; R”̈ottger et al., 2022a).
Multi-Source Training
@@
This approach often entails the amalgamation of labelled data from multiple source languages and potentially the target language in diverse ways. By synthesising information across multiple languages, this strategy is able to capture the diversities and commonalities of linguistic patterns, demonstrating significant efficacy in the cross-lingual setting. It is extensively utilised across all three transfer scenarios, serving as an augmented cross-lingual learning approach during different stages of training or fine-tuning. In multi-source training, the most prevalent way is to train or fine-tune the model on multiple source languages. This can be executed in the training phase of joint learning (R”̈ottger et al., 2022b; Muti and Barrón-Cedeño, 2022; Salaam et al., 2022; Deshpande et al., 2022; Vashistha and Zubiaga, 2021; Elouali et al., 2020; zahra El-Alami et al., 2022; Das et al., 2022; Hande et al., 2021a; Markov et al., 2021; Chiril et al., 2019) or the fine-tuning phase of cascade learning (Ranasinghe and Zampieri, 2021a; Gupta et al., 2022) in conjunction with the target language, or in isolation from the target language for zero-shot transfer (Ryzhova et al., 2022; Montariol et al., 2022). Deshpande et al. (2022) offer a more focused perspective by narrowing the linguistic scope to specific language families. By training models on languages within one language family, they then assess the model’s performance on each member of that family, revealing the interplay of linguistic knowledge within related languages. It is worth noting that merging training examples from different languages can be detrimental to cross-lingual performance in cases where offensive language domains are too distant (Glavaš et al., 2020).
Multi-Channel Learning
@@
This strategy typically refers to the use of multiple types of input representations or sources concurrently for a single task, where multiple embeddings or pre-processed versions of the text are often channelled as parallel input streams. In this architecture, each channel processes the input independently through parallel layers. These independent processing channels are eventually combined based on their outputs for subsequent learning or classification stages. Such a design ensures that diverse input representations are holistically integrated, capturing a broader spectrum of linguistic knowledge across languages. Multi-channel learning finds its applications mainly in the zero-shot transfer and joint learning scenarios. When applied to cross-lingual offensive language detection, this approach often treats datasets from different languages as distinct input channels. Each channel adopts different neural networks or PLMs based on diverse input languages, and these channels are then merged into the final classification stage (Kumar and Sachdeva, 2022; Stappen et al., 2020; Sohn and Lee, 2019). A significant improvement to this strategy is the incorporation of machine translation. By translating labelled source data to pseudo target data, low-resource target data is enriched and a bridge is constructed between source and target languages, enhancing knowledge relationships and a more seamless transfer of languages (Pamungkas and Patti, 2019; Pamungkas et al., 2020, 2021a; Jiang and Zubiaga, 2021; Ibrohim and Budi, 2019a).
Multi-Task Learning
@@
A joint learning environment training a model on several interrelated tasks simultaneously, where the proficiency learned from one task can boost performance on another. A key point of this strategy is the shared representation of hidden features across specific layers of the model. Such shared layers enable different tasks to mutually benefit from common representations, while also preserving their distinctiveness through task-specific output layers. The training process is holistic, aiming to optimise the model’s performance across all tasks. This is typically achieved by incorporating the loss functions from each task. Multi-task learning emerges as a powerful method for cross-lingual offensive language detection within the parameter-level transfer techniques across three transfer scenarios. Models can be trained on offensive language detection and a wide range of related auxiliary tasks, such as Sentiment Analysis, Named Entity Recognition (NER), Dependency Parsing, and Part-Of-Speech (POS) Tagging (Montariol et al., 2022; Hande et al., 2021a; Ousidhoum et al., 2019). Additionally, they can be fine-tuned on data in various offensive categories, such as aggressive content (Ranasinghe and Zampieri, 2020, 2021c). This multi-task training strategy enables the model to identify and process offensive content with higher accuracy and enhances the model’s ability to grasp linguistic nuances across languages.
Meta Learning

### v2

@@ 6.1.1. Label-Based Transfer
It leverages parallel corpora to transfer annotations from a source to a target language. Parallel corpora consist of sentence-aligned source-to-target data, ensuring higher quality than machine-generated translations. Since data is paired between source and target languages, labels in the source language can be directly projected onto the target language. For example, if a sentence in the source language is labelled as hateful, its corresponding text in the target language can be inferred to the same label. It can effectively create labelled data for the target language without manual annotation, especially when building datasets for languages where annotating from scratch might be challenging due to linguistic complexities or a lack of annotators.
@@ 6.2. Feature Transfer
Feature transfer transforms linguistic features to aid in HS detection. Rather than translating source texts or projecting source labels, they focus on extracting salient linguistic features from source and target languages, and aligning them into a shared feature space, retaining the essence of the text consistent across languages. Cross-lingual Word Embeddings (CLWEs), also known as Multilingual Word Embeddings (MWEs), are commonly used for feature transfer. They train monolingual word embeddings (WEs), such as Word2Vec or FastText, on multiple languages, yielding vectors that capture semantic similarities relying on shared representations between languages and thus enabling representations that are common across languages. In cross-lingual offensive language detection, various approaches utilise existing CLWEs or retrofit CLWEs. We discuss three main techniques: multilingual distributional representations, multilingual contextualised representations, and retrofitting WEs.
@@ 6.2.1. Multilingual Distributional Representations
Pre-trained distributional WEs are most commonly used as embedding layers in neural networks for different language data inputs. Among these, MUSE embeddings have emerged as a predominant choice. They are extensively utilised to extract multilingual features and are often integrated with both traditional machine learning models, such as Gradient Boosted Decision Trees (Arango et al., 2022), and advanced deep learning architectures (i.e. LSTM (Pamungkas and Patti, 2019; Pamungkas et al., 2020, 2021a; Arango et al., 2021; Burtenshaw and Kestemont, 2021), BiLSTM (Bigoulaeva et al., 2021, 2022, 2023; Jiang and Zubiaga, 2021), CNN-GRU (Deshpande et al., 2022; Aluru et al., 2021, 2020), Capsule Networks (Jiang and Zubiaga, 2021), Sluice networks (Ousidhoum et al., 2019)), and BERT (Pamungkas et al., 2020)). Other notable multilingual distributional vectors, such as Multilingual GloVe and FastText, have also been employed in architectures like BiLSTM (Chiril et al., 2019) and Capsule Networks (Jiang and Zubiaga, 2021). In addition, Babylon multilingual embeddings have been paired with Sluice networks for multilingual HS detection (Ousidhoum et al., 2019). Bansal et al. (2020) integrate bilingual switching features into the Hierarchical Attention Network (HAN) architecture, enhancing the transfer knowledge for cross-lingual detection.
@@ 6.2.3. Retrofitting Word Embeddings
Retrofitting pre-trained WEs by infusing multilingual domain knowledge stands as a promising strategy to amplify the semantic relationships across languages, thereby enhancing the efficacy and scalability of models in cross-lingual HS detection. Pant and Dadu (2020) employ a supervised FastText model trained on the sarcasm detection corpus (Swami et al., 2018) to improve the identification of HS content in English and Hindi languages. Arango et al. (2021) construct HS-specific WEs by aligning monolingual Word2Vec embeddings using hate-specific bilingual dictionaries (such as HurtLex), which are able to capture non-traditional translations of words between languages. Additionally, Hahn et al. (2021) learn semantic subspace-based representations to model profane languages on both word and sentence levels. Jiang and Zubiaga (2023) propose a cross-lingual domain-aware semantic specialisation system to construct sexism-aware WEs. They retrofit pre-trained FastText word vectors by integrating in-domain and out-of-domain linguistic knowledge (such as lexico-semantic relations) into the specialised feature space.
@@ 6.2.4. Infusing Additional Features
Additional external features are commonly used in identifying HS across languages, such as domain-specific (Pamungkas and Patti, 2019; Hahn et al., 2021), language-specific (Zhou et al., 2023; Arango et al., 2021), and typographic features (Kumar and Sachdeva, 2022; Markov et al., 2021). Specifically, domain-specific features are essential to assist the hate-related knowledge transfer. The multilingual lexicon, HurtLex, has been employed in multiple studies (Pamungkas et al., 2020, 2021a; Pamungkas and Patti, 2019; Jiang and Zubiaga, 2021). Some researchers also utilise hateful word pairs to construct domain-aware or semantic-based representations across languages (Jiang and Zubiaga, 2023; Hahn et al., 2021; Arango et al., 2021), due to the lack of hate-specific patterns in general-purpose multilingual embeddings. In addition, infusing language-specific features can help integrate linguistic and cultural knowledge for geographically sensitive tasks. It can be in different forms, such as language switching pattern matrix (Bansal et al., 2020), bilingual pairs (Jiang and Zubiaga, 2023), cross-cultural similarities (Arango et al., 2021), and social dynamics among users (Nozza, 2021). This feature can provide a richer understanding of cultural dimensions in hateful content (Zhou et al., 2023). Furthermore, typographic features provide a language-agnostic and domain-agnostic perspective to help identify online hate due to their shared meanings across languages, which includes capitals, punctuations, and emojis (Kumar and Sachdeva, 2022; Markov et al., 2021; Corazza et al., 2020). Stylometric features also remain persistent with respect to language variations, as the writing style of toxic content can be correlated to emotional profile of social media users (Markov et al., 2021).
@@ 6.3. Parameter Transfer
The performance of NLP models is dependent on model parameters. Parameter transfer assumes that prior distributions of parameters are transferred between different languages within one model or between individual models. Most parameter transfer approaches rely on inductive transfer learning in cross-lingual offensive language detection (Pan and Yang, 2010). Inspired by Pikuliak et al. (2021), we outline three distinct scenarios for surveyed studies, depending on how the source and target data are utilised during the process of parameter transfer, namely zero-shot transfer, joint learning, and cascade learning (see Figure 5). Only source data is used for training in zero-shot transfer, while both source and target data are applied during the training process in joint learning and cascade learning. Joint learning utilises both simultaneously to train the model, but cascade learning leverages them in the different stages of training.
@@
| Multilingual PLMs | ✓✓ | ✓ | ✓✓ |
| Domain Adaptation | ✓ | ✓✓ | |
| Multi-source Training | ✓✓ | ✓✓ | ✓ |
| Multi-channel Learning | ✓ | ✓✓ | |
| Multi-task Learning | ✓ | ✓✓ | ✓ |
| Meta Learning | ✓ | ✓ | |
| Ensemble Learning | ✓ | ✓✓ | ✓ |
@@ 6.3.1. Transfer Scenarios
Scenario 1: Zero-Shot Transfer. It is a promising scenario for cross-lingual offensive language detection when labelled data in the target language is scarce or lacking, by enabling detection in a language not seen during training. The model is trained on labelled source data from one or more languages not including the target language, and is then tested on the unseen target language. This is sometimes called direct or model transfer. We found 25 papers under this transfer paradigm, including subtypes:
Single Source to Single Target: Considered a type of domain adaptation (see §6.3.2), models are trained on labelled data from a single source language, then used on an unseen target language (Glavaš et al., 2020; Firmino et al., 2021; Tiţa and Zubiaga, 2021; Rodríguez et al., 2021; Pelicon et al., 2021a; Nozza, 2021; Montariol et al., 2022; Eronen et al., 2022; Zhou et al., 2023; Gupta et al., 2022; Ranasinghe and Zampieri, 2021a, b; Gaikwad et al., 2021; Muti and Barrón-Cedeño, 2022; Das et al., 2022; Plaza-del arco et al., 2023). It is the most frequently observed (62.5%) of zero-shot transfer strategies in our survey.
Multiple Sources to Single Target: Models are trained on labelled data from multiple source languages (excluding the target language) and then evaluated on a single target language (Ryzhova et al., 2022; Montariol et al., 2022; Aluru et al., 2020). It is a type of multi-source training strategy where the combination of training data excludes the target (see §6.3.2).
@@
It is worth highlighting that these methods rely on multilingual PLMs, especially often used for Single Source to Single Target and Multiple Sources to Single Target scenarios. mBERT and XLM-R are two of the most prominent multilingual PLMs in surveyed papers, because of their popularity, widespread adoption and SOTA performance across multilingual NLP tasks. They together account for 80.8% of papers in the zero-shot transfer scenario. Additionally, some papers release domain- and language-specific PLMs in a multilingual setting (Gupta et al., 2022), choose other multilingual PLMs like XLM (Stappen et al., 2020) and language-specific MuRIL (Gupta et al., 2022; Das et al., 2022), or combine MWEs with monolingual neural models like LSTM (Pamungkas and Patti, 2019; Jiang and Zubiaga, 2021) and BERT (Pamungkas et al., 2020).
Scenario 2: Joint Learning. It typically involves training a model between languages based on multiple channels, tasks or objectives simultaneously, or jointly assimilating multiple types of data or modalities. The model might have some parameters that are shared across languages as well as some that are specific to each, so it is also called parameter sharing (Pikuliak et al., 2021). The transfer of knowledge happens only on shared parameters. Changes to the shared parameters can affect each other between languages. They are interrelated and can provide complementary information.
@@
Scenario 3: Cascade Learning. In the cascade learning scenario, both source and target data are used at different times during the training. First, an existing PLM is directly used or a new model is trained with source data, and then the trained model is fine-tuned with target data, where subsets of parameters are shared between different stages of training and fine-tuning. Different strategies for training and fine-tuning have been used in 19 of our surveyed papers.
Many use existing SOTA pre-trained multilingual models serving as the foundation,171717Cascade learning fine-tunes PLMs using target data, while zero-shot transfer fine-tunes PLMs using non-target data. while some studies train a model with one or multiple source datasets. Subsequently, the fine-tuning stage in cascade learning is applied across an entire (Kapoor et al., 2019; Firmino et al., 2021; Aluru et al., 2021; Pelicon et al., 2021b; Husain and Uzuner, 2022) or only a limited subset of target data (i.e. few-shot learning) (Ranasinghe and Zampieri, 2021a; Pelicon et al., 2021b), optimising the model adaptability. Röttger et al. (2022a) indicate that an initial training phase on source data (English) could increase model performance when there is little fine-tuning data in the target language, where source data can partly substitute target data in few-shot settings. Some researchers delve into consistent fine-tuning (Ryzhova et al., 2022; Mozafari et al., 2022; Awal et al., 2023) or iterative fine-tuning (i.e. meta learning) (Mozafari et al., 2022; Awal et al., 2023), performing multiple rounds of fine-tuning on the PLMs to further enhance its ability to transfer knowledge.
Additionally, given the limited availability of target data, data augmentation techniques have emerged. They aim to enrich the target dataset to enable fine-tuning, by enhancing label diversity or expanding the text corpus. Some studies have attempted to generate pseudo labels for unlabelled target datasets (Bigoulaeva et al., 2021, 2022, 2023; Zia et al., 2022) (see §6.1.1), while others augment and refine pseudo texts by exploring advanced techniques, such as word alignment to project samples (Shi et al., 2022), adversarial algorithm to generate attack samples (Ryzhova et al., 2022), and domain-specific target data filtering (Glavaš et al., 2020).
@@ 6.3.2. Hybrid Transfer Strategies
According to the utilisation of source and target data, these strategies can be broadly covered in one or more parameter transfer scenarios mentioned in §6.3.1. That is, these strategies are not mutually exclusive. A single scenario can seamlessly integrate multiple transfer strategies (see Table 6).
@@
Multilingual Pre-Trained Language Models. These models are regarded as a cornerstone of parameter-level transfer techniques for low-resource scenarios, especially in cross-lingual offensive language detection. They are trained on large-scale multilingual corpora, learning commonalities and differences between languages. Therefore, a single PLM can share the same underlying sub-word vocabulary and semantic representations across languages (Devlin et al., 2019). Multilingual PLMs can be initialised as the base model with pre-trained parameters, then fine-tuned on available HS data in all three scenarios of parameter transfer (zero-shot transfer, joint learning and cascade learning) (Muti and Barrón-Cedeño, 2022; Firmino et al., 2021). Beyond this, they can also employed in the other two transfer levels (data and feature). They are able to augment target datasets by predicting pseudo labels for unlabelled data and generating synthetic target texts. And they can be used as multilingual representations, where subsequent models are constructed based on these source representations (Rodríguez et al., 2021).
@@
Domain Adaptation. It involves taking a model trained (or pre-trained) in the source language and adapting its parameters to initialise a new model for the low-resource language, aiming to leverage source data to improve performance on target data (Lee et al., 2021; Glavaš et al., 2020). Its advantage is that it can capture offensive language knowledge and linguistic similarities between the languages within the model’s parameters. In the zero-shot transfer scenario, parameters of the trained model are entirely or partially used on the target data for prediction (Gaikwad et al., 2021; Ranasinghe and Zampieri, 2021c). In addition, cascade learning adopts a more subtle approach. That is, the entire or specific subsets of the model’s parameters (like specific layers) are frozen to retain their original state (Ranasinghe and Zampieri, 2021a), while others are fine-tuned to better align with the target data, often using a smaller target dataset (as few-shot learning) (Pelicon et al., 2021b; Röttger et al., 2022a).
@@
Multi-Source Training. This approach often entails the amalgamation of labelled data from multiple source languages and potentially the target language in diverse ways. By synthesising information across multiple languages, this strategy is able to capture the diversities and commonalities of linguistic patterns, demonstrating significant efficacy in the cross-lingual setting. It is extensively utilised across all three transfer scenarios, serving as an augmented cross-lingual learning approach during different stages of training or fine-tuning. In multi-source training, the most prevalent way is to train or fine-tune the model on multiple source languages. This can be executed in the training phase of joint learning (Röttger et al., 2022b; Muti and Barrón-Cedeño, 2022; Salaam et al., 2022; Deshpande et al., 2022; Vashistha and Zubiaga, 2021; Elouali et al., 2020; zahra El-Alami et al., 2022; Das et al., 2022; Hande et al., 2021a; Markov et al., 2021; Chiril et al., 2019) or the fine-tuning phase of cascade learning (Ranasinghe and Zampieri, 2021a; Gupta et al., 2022) in conjunction with the target language, or in isolation from the target language for zero-shot transfer (Ryzhova et al., 2022; Montariol et al., 2022). Deshpande et al. (2022) offer a more focused perspective by narrowing the linguistic scope to specific language families. By training models on languages within one language family, they then assess the model’s performance on each member of that family, revealing the interplay of linguistic knowledge within related languages. It is worth noting that merging training examples from different languages can be detrimental to cross-lingual performance in cases where offensive language domains are too distant (Glavaš et al., 2020).
@@
Multi-Channel Learning. This strategy typically refers to the use of multiple types of input representations or sources concurrently for a single task, where multiple embeddings or pre-processed versions of the text are often channelled as parallel input streams. In this architecture, each channel processes the input independently through parallel layers. These independent processing channels are eventually combined based on their outputs for subsequent learning or classification stages. Such a design ensures that diverse input representations are holistically integrated, capturing a broader spectrum of linguistic knowledge across languages. Multi-channel learning finds its applications mainly in the zero-shot transfer and joint learning scenarios. When applied to cross-lingual offensive language detection, this approach often treats datasets from different languages as distinct input channels. Each channel adopts different neural networks or PLMs based on diverse input languages, and these channels are then merged into the final classification stage (Kumar and Sachdeva, 2022; Stappen et al., 2020; Sohn and Lee, 2019). A significant improvement to this strategy is the incorporation of machine translation. By translating labelled source data to pseudo target data, low-resource target data is enriched and a bridge is constructed between source and target languages, enhancing knowledge relationships and a more seamless transfer of languages (Pamungkas and Patti, 2019; Pamungkas et al., 2020, 2021a; Jiang and Zubiaga, 2021; Ibrohim and Budi, 2019a).
@@
Multi-Task Learning. A joint learning environment training a model on several interrelated tasks simultaneously, where the proficiency learned from one task can boost performance on another. A key point of this strategy is the shared representation of hidden features across specific layers of the model. Such shared layers enable different tasks to mutually benefit from common representations, while also preserving their distinctiveness through task-specific output layers. The training process is holistic, aiming to optimise the model’s performance across all tasks. This is typically achieved by incorporating the loss functions from each task. Multi-task learning emerges as a powerful method for cross-lingual offensive language detection within the parameter-level transfer techniques across three transfer scenarios. Models can be trained on offensive language detection and a wide range of related auxiliary tasks, such as Sentiment Analysis, Named Entity Recognition (NER), Dependency Parsing, and Part-Of-Speech (POS) Tagging (Montariol et al., 2022; Hande et al., 2021a; Ousidhoum et al., 2019). Additionally, they can be fine-tuned on data in various offensive categories, such as aggressive content (Ranasinghe and Zampieri, 2020, 2021c). This multi-task training strategy enables the model to identify and process offensive content with higher accuracy and enhances the model’s ability to grasp linguistic nuances across languages.